OpenStudy (nilankshi):

five year hence,the age of jacob will be three times that of his son.five years ago.jacob's age was seven times that of his son.what is thair present age?

OpenStudy (aravindg):

simple first write the qns using x and y

OpenStudy (anonymous):

Ages x,y x/y = 7 (x+10)/(y+10) = 3 Solve simultaneously